PIP-II report for May 25, 2018

Highlights:

  • Last week of PIP2IT operations – Maximum parameters achieved: 5mA, 24 msec at 20 Hz.
  • First collaboration meeting with INP Orsay on SSR2 cavity development.
  • Three SSR1 cavities with high power coupler are ready for string assembly.
  • Site Prep Coordination Meeting was held.

 

Project Management (Marc Kaducak):

  • We are reviewing resumes for the Procurement Manager position and have assembled an interview panel in conjunction with the Procurement Department.
  • Worked with Accelerator division L2 Managers to identify critical resources needed in their WBS areas.
  • Project controls team made rapid progress in updating the schedule for SSR1 and SSR2. OPSS reviewed the Safety Systems area of the resource loaded schedule and we are incorporating their recommendations.

 

Technical Integration (Arkadiy Klebaner):

  • Advancing the Interface Control Documents development – 60% completion.
  • Issued an example document list for design deliverables.
  • Working towards an updated list of requirements for PIP2IT FRS (ED0001223), to align with PIP-II pulsed beam mode.
  • Distributed OPA guidance on spares policy, which includes discussion of types of spares, and asked L3 managers to put together a list for their respective systems.
  • Held first design change request (DCR) review.
  • Organized discussions between Accelerator Systems and Accelerator Integration managers to help identify requirements, interfaces, and design deliverables.

 

International Coordination (Steve Holmes & Lia Merminga):

  • DAE Labs:
    • RFPI TRS rev 8 released for final review and approval.
    • Fermilab-DAE teleconference to discuss status/plan for supporting LLRF systems at RRCAT/HTS-2 and Fermilab/PIP2IT
  • INP Orsay:
    • First collaboration meeting with Orsay colleagues about SSR2 cavity.

 

SRF & Cryo Systems (Genfa Wu):

HWR

  • HWR CM Production
    • FNAL and ANL staff are presently reviewing the HWR CM assembly and testing plans and procedures with the goal of optimizing the CM delivery date to Fermilab. A meeting is planned on 6/1/2018 to discuss the potential revisions of the schedule.
  • HWR Integrated Cold Test with Power Coupler
    • The RF system at ANL that supports HWR coupler and cavity testing is being upgraded with assistance from FNAL personnel to utilize a higher power circulator and improved interlocks.

SSR1

  • SSR1 Cavities, coupler and tuners
    • S107 test with high power coupler was completed. The Q0 was below specification at 10 MV/m. However, this cavity can still be accepted into a string assembly. Several cavities in a SSR1 cryomodule requires to be operated at lower gradient around 4-6 MV/m.
  • SSR1 Cryomodule
    • Three cavities with high power coupler are ready for string assembly. S1H-NR-112, S1H-NR-113 and now with S1H-NR-107.
    • Current leads prototype design is completed after incorporating lessons-learned from HWR current leads vacuum leak. Parts are being procured for this prototype..

SSR2

  • A first collaboration meeting with Orsay colleagues about SSR2 was productive. Some preliminary design collaboration can commence.

HB650

  • Package for procurement of Nb for prototype cavities (qty 2) being prepared. Requisition to be entered in the near future. Material specification is being revised. A trip to JLAB will help finalize the material specification by learning lessons from LCLS-II material ordering.
  • 5-cell beta=0.90 cavity B9A-AES-010 dressing mock up is in progress.
  • Coupler clean room tooling design is completed. Parts will be ordered soon. The tooling is for the new design of the coupler which we have received. The goal is to develop and demonstrate the particle free assembly procedure.

Cryogenics Plant

  • RLS development for the Cryoplant is in progress.
  • Collaboration visit to ESS was completed. Useful information was obtained in cryogenic plant design and operation.

Cryogenics Distribution System

  • RLS development for the CDS is in progress.
  • Collaboration visit to ESS was completed. Useful information was obtained in cryogenic distribution system design and implementation.
  • Met with Conventional Facilities manager to discuss location of CDS transfer line trench/penetration into the Linac tunnel and general transfer line routing methods.

 

Accelerator Systems (Paul Derwent):

High Power RF

  • The new 162.5 MHz RF distribution design for HWR cavity and input coupler testing was commissioned.
  • The remaining HWR power amplifiers contract has been awarded, delivery will be in stages starting in Dec ‘18 and concluding in March ’19.
  • The purchase requisition for 6-1/8″ coax components for Coupler Test at MDB has been submitted in the AD Req. system.
  • The 162.5 MHz 3 kW circulator, which failed during a coupler test at Argonne, has been tested on the bench.  The circulator will be returned to the vendor for analysis.

Vacuum

  • Presented the analysis of RFQ vacuum burst before/after the modification of pumping port plate at the right side of module #3.
  • Planned the reconfiguration of last section PIP2IT/MEBT to further test the performance of DPI-Fast Valve with better precision, and to have a preliminary baseline of vacuum quality as a part of preparation for HWR.
  • Prepared the drawing package for 200 Ohm Kicker production.

 

Linac Installation & Commissioning (Fernanda G. Garcia):

PIP2IT:

  • This week marked the last week of operations in PIP2IT.
    • Achieved maximum of 5mA, 24 msec at 20 Hz to the SNS beam dump.
    • Long run attempted at 14ms, 5mA at 20 Hz.
  • Hosted tours at the facility.
  • Long shutdown for Cryogenic Transfer Line installation and cave reconfiguration for cryomodules installation starts next week, May 29th.
  • Internal review of the 200-ohm kicker design has been postponed to second half of July.
  • Group continues working on RLS and engineering documentation.

Test Infrastructure:

  • PIP2IT Cryogenic Transfer Line
    • The factory acceptance test did not pass the final QA test and will be delayed a few weeks.
    • However, PIP2IT operation run is complete and on May 29th activities associated with the cave extension and preparation for the cryogenic transfer line installation and HWR arrival will commence as scheduled.
  • Continue working on RLS and documentation.

Installation:

  • Discussed ICD issues with CF and systems engineering.
  • TD design group has begun creating placeholder models for PIP-II magnets.
  • Received preliminary approval from MSD to use A0 high bay North End as a PIP-II staging area during installation. This does not meet all expected storage and logistics needs but, is a good start.

 

Conventional Facilities (Steve Dixon):

  • Completed field work for the soil borings portion of the geotechnical investigation report.
  • Met with Interlocks group to understand space and infrastructure requirements.
  • Site Prep Coordination Meeting – Met with the design team to coordinate progress for the Site Prep package.
  • Began updating Basis of Estimate documents with FY18 cost estimates.